# Counting with the quantum alternating operator ansatz

### v1

@@ 5 Solving #P-Complete Problems with VQCount
Figure 9: Number of post-selected samples needed to achieve an approximate count with error tolerance $\varepsilon$ with depth $p=3$ QAOA circuits for 1-in-3SAT instances at $\alpha=2/3$. The inset panel shows the scaling behavior with the inverse of the error tolerance $\varepsilon$ for $n=24$ variables, for which a polynomial fit gives $O(\varepsilon^{-0.79})$. Solid lines represent the mean, and shaded regions show the standard error of the mean.
6 Conclusion
This work introduces VQCount, a sampling-based variational quantum heuristic for approximate model counting based on QAOA and the JVV theorem. We establish the soundness of VQCount in a limit where it is shown to reproduce Grover’s algorithm. We show that, compared to previous work [17], VQCount offers an exponential improvement in the number of samples required for obtaining an estimate within a multiplicative factor of the exact count whenever the number of solutions is exponential. We then investigate the performance of VQCount as a heuristic numerically, using tensor network contraction to simulate the quantum circuit component of the algorithm. For the counting problems ##\##NAE3SAT and ##\##1-in-3SAT, we establish trade-offs between solution sampling success rate and sampling uniformity. The systematic improvement of performance as a function of circuit depth (Fig. 9(a)) suggests that VQCount with circuits of (potentially variable) larger depth than those amenable to classical simulation with reasonable resources may eventually become competitive with classical heuristics for approximate counting. In this work, the only post-processing used is postselection to reject samples outside the solution set. However, more elaborate post-processing, such as rejection-free cluster updates at zero temperature [48], could be employed to alleviate nonuniformity and reduce postselection in the sampling.
7 Data and code availability
Data is available upon reasonable request, and code can be found on GitHub.
8 Acknowledgments
@@ Appendix A Determining the minimum number of samples required to reach a given accuracy with VQCount
To determine the minimum number of samples necessary to reach an approximate count with a multiplicative error $\varepsilon$, the VQCount algorithm was run by increasing the number of solutions sampled from the optimized QAOA circuit until the approximate number of solutions was inside the given error bounds. Fig. 10 shows the count accuracy, i.e., the approximate number of solutions over the exact number of solutions, for given error bounds. VQCount systematically underestimates the number of solutions, as seen most prominently in the ##\##NAE3SAT curve for $\alpha=1$.

### v2

@@ V Solving #P-Complete Problems with VQCount
Figure 9: Number of post-selected samples needed to achieve an approximate count with error tolerance $\varepsilon$ with depth $p=3$ QAOA circuits for 1-in-3SAT instances at $\alpha=2/3$. The inset panel shows the scaling behavior with the inverse of the error tolerance $\varepsilon$ for $n=24$ variables, for which a polynomial fit indicates an asymptotic scaling of $\varepsilon^{-0.78\pm 0.01}$. Solid lines represent the mean, and shaded regions show the standard error of the mean.
Classical sampling-based approximate counting methods struggle to generate samples almost uniformly from complex solution spaces, prompting the use of hashing-based counting algorithms. VQCount proposes a tunable tradeoff between classical and quantum resources (postselection cost vs. ansatz depth) for addressing the solution sampling challenge.
VI Conclusion
This work introduces VQCount, a sampling-based variational quantum heuristic for approximate model counting based on QAOA and the JVV theorem. We establish the soundness of VQCount in a limit where it is shown to reproduce Grover’s algorithm. We show that, compared to previous work [52], VQCount achieves an exponential improvement in the number of samples required for obtaining an estimate within an arbitrarily small multiplicative factor of the exact count whenever the number of solutions is exponentially large.
We then numerically investigate the performance of VQCount, with both GM-QAOA and regular QAOA as sampling subroutines, as a heuristic (without theoretical guarantees), using tensor network contraction to simulate the quantum circuit component of the algorithm. For the counting problems #\#NAE3SAT and #\#1-in-3SAT, we establish trade-offs between solution sampling success rate and sampling uniformity. These results establish VQCount as a viable framework for approximate counting using VQAs.
The implementation introduced here can serve as a baseline for future improvements in VQCount, including problem-aware mixers, warm-start strategies, and improved parameter optimization heuristics. While the current scaling of VQCount remains worse than that of modern classical solvers, Fig. 8(a) indicates consistent performance improvement with increasing ansatz depth. This improvement leaves open the possibility that VQCount with circuits of (potentially variable) larger depth may eventually become competitive with classical heuristics for approximate counting.
Prior work has shown that accurate approximate counting is possible even with non-uniform sampling [24]. It is plausible that similar arguments may also hold for VQCount. Finally, the only post-processing used in this work is postselection to reject samples outside the solution set. However, more elaborate post-processing, such as rejection-free cluster updates at zero temperature [43], could be employed to reduce the postselection cost.
Acknowledgments
@@
Data availability statement
The data that support the findings of this study are openly available at the following URL/DOI: https://github.com/QuICoPhy-Lab/VQCount [17].
@@ Appendix A Determining the minimum number of samples required to reach a given accuracy with VQCount
To determine the minimum number of samples necessary to reach an approximate count with a multiplicative error $\varepsilon$, the VQCount algorithm was run by increasing the number of solutions sampled from the optimized QAOA circuit until the approximate number of solutions was inside the given error bounds. Note that the minimum number of samples includes those discarded by postselection. Fig. 10 shows the count accuracy, i.e., the approximate number of solutions over the exact number of solutions, for given error bounds. VQCount systematically underestimates the number of solutions, as seen most prominently in the #\#NAE3SAT curve for $\alpha=1$.